OpenStudy (anonymous):

For the geometric series 1 + 3 + 9 + 27 + . . . , find the sum of the first 10 terms

OpenStudy (anonymous):

a=1 r=3/1=3 \[S10=\frac{ a \left( 1-r ^{10} \right) }{1-r }\] calculate

OpenStudy (anonymous):

A. 19,683 B. 29,523 C. 29,524 D. 88,573

OpenStudy (uri):

1+3+9+27+81+243+729+2187+6561+19683=29,524 I know there is a formula..but i didn't know it so that i did....multiply its basically 3 times table...Like 1 x 3=3. 3 x 3 =9 9 x 3=27 27 x 3 81 and etc

OpenStudy (anonymous):

thanks

OpenStudy (uri):

yw! :)